OpenStudy (anonymous):

2. Use the method of cylindrical shells to find the volume V generated by rotating the region bounded by the given curves about y = -5. y=x^2, x=y^2 Volume =

OpenStudy (amistre64):

y = -5. y=x^2, x=y^2

OpenStudy (amistre64):



OpenStudy (amistre64):

i take it they meet at 0 and 1?

OpenStudy (amistre64):